OpenStudy (gdgutierrez):

Help... Find the slope of tangent line x(t)= sqrt(t), y(t) = (1/4)(t^2 -4), t>0 Answer: 8 I have no clue....

ganeshie8 (ganeshie8):

are u given a "point" at which u want the slope of tangent line ?

OpenStudy (gdgutierrez):

no. That is literally the entire question...

OpenStudy (anonymous):

first calculate dy/dt and dx/dt now calculate dy/dx by (dy/dt)/(dx/dt) this is the slope

OpenStudy (anonymous):

or eliminate t and find y as a function of x then differentiate it.

OpenStudy (anonymous):

i like that way if you are able to do it

OpenStudy (gdgutierrez):

But how is the answer 8??

OpenStudy (anonymous):

no it;s function of x or t for any number you have to know specific point where tangent asked.

OpenStudy (gdgutierrez):

so if I eliminate t between x and y I would get the same answer as if I tool dy/dx right?

OpenStudy (gdgutierrez):

I feel like I am missing something because elimination gets me x^3 and dy/dx gets me t(sqrt(t))

OpenStudy (anonymous):

after eliminate t you gate equation of curve as y= f(x) now you have to find dy/dx.

ganeshie8 (ganeshie8):

tsqrt(t) = x^2sqrt(x^2) = x^2.x = x^3

ganeshie8 (ganeshie8):

so nthng to confuse okay

OpenStudy (gdgutierrez):

so now what's missing is the point huh?

ganeshie8 (ganeshie8):

obviously ! slope is not a constant, so u wont get 8 as answer unless u r a given a point

OpenStudy (anonymous):

x(t)= sqrt(t), y(t) = (1/4)(t^2 -4), square first you get t= x^2 now put this value of t in second you will get y= function of x

OpenStudy (gdgutierrez):

My professor is too old then to forget to put a point then. Thanks guys. Appreciate it.

ganeshie8 (ganeshie8):

lol thats funny ;)

ganeshie8 (ganeshie8):

u wlc !

OpenStudy (anonymous):

A point technically is not needed to find the slope of the mutual tangent lines.

OpenStudy (gdgutierrez):

do you know how to get 8 then?